# Restricted Path Characteristic Function Determines the Law of Stochastic Processes

### v2

@@ 5.1. Separation of points
Let $\gamma_{1}$, $\gamma_{2}\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$ be paths on $\mathbb{R}^{d}$ of bounded $p$-variations; $21\leq p<$ 2. Suppose that their signatures differ at the $n^{\text{th}}$ level, namely that $\pi_{n}[S(\gamma_{1})]\neq\pi_{n}[S(\gamma_{2})]$. Then, for any $k\geq n+1$ one can find $M\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ such that $\mathscr{D}_{\gamma_{1}}(M)\neq\mathscr{D}_{\gamma_{2}}(M)$.
Chevyrev–Lyons [7, Theorem 4.8] asserts that for any $X\in{\mathscr{E}}^{\infty}$ (which, by definition, is a subset of $\widehat{{\mathscr{E}}}$) with $\pi_{n}(X)\neq 0$, one can find $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{sp}(m)\right)$ for some $m\geq\max\left\{2,\frac{n}{3}\right\}$ such that $\Phi_{X}(M)\neq 0$. In contrast, Corollary 5.1 establishes that linear mappings taking values in the smaller matrix algebra $\mathfrak{t}(k,\mathbb{R})$ separate points over a larger subset of the tensor algebra containing the expected signatures.
5.2. Uniqueness of restricted path characteristic function
Motivated by Corollary 5.1, let us consider the restriction of the development map $\mathscr{D}_{\gamma}$ to the domain $\mathcal{L}\left(\mathbb{R}^{d};\bigoplus_{k=0}^{\infty}\mathfrak{t}(k,\mathbb{R})\right)$. This serves as the foundation for our novel applications of $\mathscr{D}_{\gamma}$ to hypothesis testing in §6 below.
@@ 5.2. Uniqueness of restricted path characteristic function
Let $21\leq p<$ 2 and $\gamma\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$. Let $\mathscr{D}_{\gamma}:\mathcal{L}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)\to G$ be the unitary development, namely that the Cartan development defined in (5) with $\mathfrak{g}=\bigoplus_{k=0}^{\infty}\mathfrak{su}(k)$. We call the restriction of $\mathscr{D}_{\gamma}$ to $\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}\right)$ the restricted unitary feature of $\gamma$. Its expectation is called the restricted path characteristic function (restricted PCF) of $\gamma$.
We establish the uniqueness of restricted PCF under the condition for the decay rate of expected signature by applying the uniqueness of expected signature ([7, Proposition 6.1] ):
Theorem 5.3.
Let $\gamma_{1},\gamma_{2}$ be $\mathcal{V}_{1}\left([0,T],\mathbb{R}^{d}\right)$-valued stochastic processes whose expected signatures have infinite radii of convergence. $S(\gamma_{1}){\buildrel d\over{=}}S(\gamma_{2})$ if and only if they have the same restricted PCF:
$\mathbb{E}\left[\mathcal{D}_{\gamma_{1}}(M)\right]=\mathbb{E}\left[\mathcal{D}_{\gamma_{2}}(M)\right]\qquad\text{for any }M\in{\mathcal{L}}\left(\mathbb{R}^{d};\bigoplus_{k=0}^{\infty}\mathfrak{t}(k,\mathbb{R})\right).$ (30)
Proof.
The expected signature has infinite radius of convergence ([7]), so it determines the law of signature. By Theorem 1.8, it is determined in turn by the restricted PCF. ∎
As an illustrative example, we apply our Main Theorem 1.8 to recover the expected signature of the $d$-dimensional Brownian motion (BM) from its expected unitary development.
Example 5.4 (Recovering the expected signature of BM).
Denote the BM on time interval $[0,T]$ as $B_{[0,T]}=\left(B^{1}_{[0,T]},B^{2}_{[0,T]},\cdots,B^{d}_{[0,T]}\right)$. The expected signature of $B_{[0,T]}$ is given in [22]:
$\mathbb{E}\left[S(B_{[0,T]})\right]=\exp\left(\frac{T}{2}\sum_{i=1}^{d}e_{i}\otimes e_{i}\right).$ (31)
It is defined via iterated integrals in the sense of Stratonovich, a.k.a. the canonical Brownian rough path [17]. Then, by definition, the generating function of $B_{[0,T]}$ is
$\Phi_{B}(M_{(\theta)}):=\widetilde{M}_{(\theta)}\Big{[}\mathbb{E}(S(B_{[0,T]}))\Big{]}=\exp\left\{\frac{T}{2}\sum_{i=1}^{d}\left[M_{(\theta)}(e_{i})\right]^{2}\right\}.$
Let $W=(w_{1}w_{2}\cdots w_{L})$ be any fixed word of length $L$. We aim at determining the $W$-component of BM’s expected signature, $\mathbb{E}[S(B_{[0,T]})]_{W}$. For notational simplicity, write $M_{i}^{W}(e_{i})$ as $M_{i}$ for $i\in\{1,2,\ldots,d\}$. Differentiate the generating function $\Phi_{B}$ to obtain that
$\frac{\partial^{L}\big{[}\Phi_{B}(M_{(\theta)})\big{]}}{\partial\theta_{w_{1}}\partial\theta_{w_{2}}\cdots\partial\theta_{w_{L}}}\bigg{|}_{\theta=0}=\begin{cases}0\qquad\text{if $r^{W}(i)$ is odd for some $1\leq i\leq d$},\\ \frac{\mathcal{C}_{W}}{\left(\frac{L}{2}\right)!}\left(\frac{T}{2}\right)^{\frac{L}{2}}\sum_{I\in\widetilde{\Sigma}(W)}\left(M_{i_{1}}M_{i_{2}}\cdots M_{i_{L}}\right)\qquad\text{if else},\end{cases}$ (32)
where
$\widetilde{\Sigma}(W):=\left\{I\in\widetilde{\sigma}(W):\,I_{p}=I_{p+1}\text{ for each odd $p\in\{1,\ldots,L\}$}\right\};$
$\widetilde{\sigma}(W):=\left\{\left(w_{\tau(1)}w_{\tau(2)}\cdots w_{\tau(L)}\right):\,\tau\in\Sigma_{L}\right\}.$
For example, $\widetilde{\Sigma}((1212))=\{(1122),(2211)\}$. As a result, the expected signature computed via Theorem 1.8 is
$\widehat{\mathbb{E}}\left[S\left(B_{[0,T]}\right)\right]_{W}=\begin{cases}\frac{1}{\left(\frac{L}{2}\right)!}\left(\frac{T}{2}\right)^{\frac{L}{2}},&\text{ if }r^{W}(i)\text{ is even and }W\in\widetilde{\Sigma}(W)\text{ for each }i\in\{1,2,\ldots,d\},\\ 0,&\text{ otherwise},\end{cases}$
which coincides with the known formula (31).
5.3. Uniqueness of orthogonal path characteristic function
As a variant of the restricted PCF in Definition 5.2, we introduce the following:
Definition 5.5.
Let $21\leq p<$ 2 and $\gamma\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$. Let $\mathscr{D}_{\gamma}:\mathcal{L}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)\to G$ be the unitary development, namely that the Cartan development defined in (5) with $\mathfrak{g}=\bigoplus_{k=0}^{\infty}\mathfrak{su}(k)$. We call the restriction of $\mathscr{D}_{\gamma}$ to $\mathcal{L}\left(\mathbb{R}^{d};\bigoplus_{k=0}^{\infty}\mathfrak{so}(k,\mathbb{R})\right)$ the orthogonal feature of $\gamma$. Its expectation is called the orthogonal path characteristic function (orthogonal PCF) of $\gamma$.
The special orthogonal matrix algebra ${\mathfrak{so}}(k,\mathbb{R})$ consists of anti-symmetric real matrices. Note for each $k\in\mathbb{N}$ that
$\mathfrak{t}(k,\mathbb{R})\subset{\mathfrak{so}}(k,\mathbb{R})\subset{\mathfrak{su}}(k).$
Without any additional regularity assumptions on the expected signature, we have:
Theorem 5.6.
Let $\gamma_{1},\gamma_{2}$ be $\mathcal{V}_{1}\left([0,T],\mathbb{R}^{d}\right)$-valued stochastic processes. Then $S(\gamma_{1}){\buildrel d\over{=}}S(\gamma_{2})$ if and only if their OPCF are the same, i.e.,
$\mathbb{E}\left[\mathcal{D}_{\gamma_{1}}(M)\right]=\mathbb{E}\left[\mathcal{D}_{\gamma_{2}}(M)\right]\qquad\text{for any }M\in{\mathcal{L}}\left(\mathbb{R}^{d};\bigoplus_{k=0}^{\infty}\mathfrak{so}(k,\mathbb{R})\right).$ (33)
Proof.
By Corollary 1.9 and $\mathfrak{t}(k,\mathbb{R})\subset{\mathfrak{so}}(k,\mathbb{R})$, the orthogonal feature separates points over tensors of positive radii of convergence. We claim that $L\circ\mathcal{D}_{\gamma}(M)$ where $L$ is a linear functional and $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\bigoplus_{k=1}^{\infty}{\mathfrak{so}}(k)\right)$ forms an algebra; i.e., for any $(L_{1},M_{1})$ and $(L_{2},M_{2})$ there exists $(L_{3},M_{3})$ such that $\Big{(}L_{1}\circ\widetilde{M_{1}}\Big{)}\bullet\Big{(}L_{2}\circ\widetilde{M_{2}}\Big{)}=L_{3}\circ\widetilde{M_{3}}$. Say $M_{i}\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{so}}(k_{i})\right)$ for $i\in\{1,2,3\}$.
To see this claim, simply note that
$\widetilde{M_{1}}(S(\gamma))\otimes\widetilde{M_{2}}(S(\gamma))=\widetilde{M_{3}}(S(\gamma))$
for
$M_{3}=M_{1}\otimes I_{k_{2}}+I_{k_{1}}\otimes M_{2},$
and that $\mathcal{D}_{\gamma}(M)$ equals the constant $1$ for $k=1$ and $M=0$. Thus, by Stone–Weierstrass [7, Lemma 4.2], the characteristic function defined in (12) with $M$ restricted to $\bigoplus_{k=1}^{\infty}{\mathfrak{so}}(k)$ determines the Borel measures on $\mathcal{G}\left(\mathbb{R}^{d}\right)$. ∎
6. Numerical experiments on hypothesis testing
6.1. A new distance for stochastic processes
According to Theorem 5.3, under the regularity condition therein, the restricted PCF of a stochastic process determines its law (and vice versa). This motivates the introduction of a new distance function for probability measures on the path space, termed as the restricted path characteristic function distance (RPCFD), by restricting the unitary development/feature to the sparse subspace $\mathfrak{t}(k,\mathbb{R})$. Similarly, one can define a orthogonal PCF-based distance function. These natural variants of the path characteristic function distance (PCFD) proposed in [15] demonstrate significant computational benefits due to the sparsity of $\bigoplus_{k=0}^{\infty}\mathfrak{t}(k;\mathbb{R})$ and $\bigoplus_{k=0}^{\infty}\mathfrak{so}(k;\mathbb{R})$.
Denote the Hilbert–Schmidt norm on $\mathbb{C}^{m\times m}=\mathfrak{gl}(m,\mathbb{C})$ as $\|\cdot\|_{\rm HS}$:
$||A||_{\rm HS}:=\sqrt{\operatorname{tr}(AA^{\ast})}.$
In the sequel we shall focus on the methodology of learning the RPCFD.
In view of Hambly–Lyons [10] and Theorem 5.3, the restricted unitary feature of a path of bounded variation only determines the path modulo tree-like equivalence. However, tree-like equivalence is mod out by lifting to time-augmented paths:
$\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right):=\left\{\tilde{X}:\,\tilde{X}(t)=(t,X(t))\text{ for }X\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)\right\}.$
Here $\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$ denotes the space of bounded $p$-variation paths into $\mathbb{R}^{d}$. We also set
$\mathcal{P}_{0}\left({\tilde{\mathcal{V}}_{1}([0,T],\mathbb{R}^{d})}\right):=\Bigg{\{}\text{Borel probability measures on $\tilde{\mathcal{V}}_{1}\left([0,T],\mathbb{R}^{d}\right)$ }$
$\qquad\qquad\text{under which the corresponding expected signature has ${\rm ROC}=\infty$}\Bigg{\}}.$ (34)
Then we introduce:
Definition 6.1.
@@ 6.1. A new distance for stochastic processes
Fix $k\geq 1$ and any Borel probability measure $\mathcal{M}$ on $\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ with full support. For $\mu,\nu\in\mathcal{P}_{0}\left({\tilde{\mathcal{V}}_{1}\left([0,T],\mathbb{R}^{d}\right)}\right)$, their restricted path characteristic function distance (RPCFD) with respect to $\mathcal{M}$ is defined as
${\rm RPCFD}_{\mathcal{M}}^{2}(\mu,\nu):=\mathbb{E}_{M\sim\mathcal{M}}\left[\big{|}\big{|}\Phi_{\mu}(M)-\Phi_{\nu}(M)\big{|}\big{|}_{\rm HS}^{2}\right],$ (35)
with the characteristic functions $\Phi_{\mu}(M):=\mathbb{E}_{X\sim\mu}\left[\widetilde{M}(S(X))\right]$ and $\Phi_{\nu}(M):=\mathbb{E}_{Y\sim\nu}\left[\widetilde{M}(S(Y))\right]$.
As $\mathfrak{t}(k,\mathbb{R})\subset{\mathfrak{su}}(k)$, RPCFD can be viewed as a restricted version of PCFD that inherits various nice properties from PCFD, including the uniform boundedness and differentiability with respect to model parameters. Moreover, since
@@
RPCFD offers a significant dimension reduction benefit (i.e., in need of fewer parameters) over PCFD while retaining the characteristic property.
@@
Likewise, based on Theorem 5.6 one may define the orthogonal path characteristic function distance (OPCFD) for stochastic processes by restricting the domain to $\bigoplus_{k=0}^{\infty}{\mathfrak{so}}(k)$. As $dim({\mathfrak{so}}(k))=k(k-1)/2$, RPCFD reduces the dimension more significantly than OPCFD, so we focus more on RPCFD. Nonetheless, most of the analysis and numerical methods below also apply to OPCFD. See Appendix B for numerical experiments of OPCFD.
To enhance the discriminative power of RPCFD on given path distributions, we propose the following trainable distance:
$\sup_{\mathcal{M}\in\mathcal{P}(\mathcal{L}(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})))}\text{RPCFD}^{2}_{\mathcal{M}}(\mu,\nu).$ (36)
@@
$\mathcal{M}\approx\frac{1}{K}\sum_{i=1}^{K}\delta_{M_{i}},$ (37)
where $M_{i}\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$, $i\in\{1,\cdots,K\}$, for some $K\geq 1$.
Note that the linear mappings $\{M_{i}\}_{1\leq i\leq K}$ can be fully parameterised. Therefore, the task of learning the optimal distribution $\mathcal{M}$ in (36) is transformed to learning the optimal parameters in the linear mappings $\{M_{i}\}_{1\leq i\leq K}$ which attain the supremum of the “empirical RPCFD”:
$\sup_{\{M_{i}\}_{1\leq i\leq K}}\frac{1}{K}\sum_{i=1}^{K}\text{RPCFD}^{2}_{\delta_{M_{i}}}(\mu,\nu):=\sup_{\{M_{i}\}_{1\leq i\leq K}}\frac{1}{K}\sum_{i=1}^{K}||\Phi_{\mu}(M_{i})-\Phi_{\nu}(M_{i})||_{\rm HS}^{2}.$ (38)

### v3

@@ 5.2. Determination of measures on path spaces
For any $\gamma_{1}\neq\gamma_{2}$ in $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$, there exists $k\geq 1$ and $M_{0}\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ such that $\mathscr{D}_{\gamma_{1}}(M_{0})\neq\mathscr{D}_{\gamma_{2}}(M_{0})$. Specifically, for any $n\geq 1$ such that $\pi_{n}\big{[}S(\gamma_{1})\big{]}\neq\pi_{n}\big{[}S(\gamma_{2})\big{]}$, we can find one such $M_{0}$ in $\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(n+1,\mathbb{R})\right)$.
For ease of notation, write
[…]
Here $\mathfrak{g}$ is a matrix Lie algebra graded by $k\in\mathbb{N}$ such that $\mathfrak{g}(k)\subset\mathfrak{gl}(k;\mathbb{C})$. In this work, we are interested in the following three choices of $\mathfrak{g}(k)$:
$\mathfrak{t}(k,\mathbb{R})\subset\mathfrak{so}(k)\subset\mathfrak{su}(k).$
As in §5.1 above, we equip $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$ with any Hausdorff topology such that for any $M\in\bigoplus_{k\geq 1}\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{su}(k)\right)$, the unitary path development $\mathscr{D}_{\cdot}(M):\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)\rightarrow\bigoplus_{k\geq 1}\mathcal{U}(k)$ is continuous. One valid choice is to equip $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$ with the $p$-variation norm [18], which makes it an inseparable Banach space. The continuity of $\mathscr{D}_{\cdot}(M)$ holds because the development of a path is the unique solution of a linear differential equation driven by the path; see Remark 1.1. The same argument holds if we replace $\mathfrak{su}(k)$ and $\mathcal{U}(k)$ by $\mathfrak{so}(k)$ and $\mathcal{O}(k)$, respectively.
Note that for $\mathfrak{g}\in\{\mathfrak{su},\mathfrak{so}\}$, $\mathscr{D}^{[\mathfrak{g}]}_{\cdot}$ is a set of bounded continuous functions on $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$ that contains the constants. In addition, for any $M_{1}\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{g}(k_{1})\right)$ and $M_{2}\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{g}(k_{2})\right)$,
$\mathscr{D}^{[\mathfrak{g}]}_{\cdot}(M_{1})\otimes\mathscr{D}^{[\mathfrak{g}]}_{\cdot}(M_{2})=\mathscr{D}^{[\mathfrak{g}]}_{\cdot}(M_{3})\qquad\text{ on $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$},$ (37)
where $M_{3}=M_{1}\otimes I_{k_{2}}+I_{k_{1}}\otimes M_{2}\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{g}(k_{1}k_{2})\right)$. Therefore, from the Stone-Weierstrass Theorem (Lemma 5.1) we deduce that
Proposition 5.5.
Continuous $\mathbb{R}$-valued functions on any compact subspace of $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$ can be uniformly approximated by path developments $\mathscr{D}_{\cdot}^{\mathfrak{g}}\,(\mathfrak{g}\in\{\mathfrak{su},\mathfrak{so}\})$.
Meanwhile, let $\mu$ and $\nu$ be Radon measures on $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$ with $1\leq p<2$. If every function in $\mathscr{D}_{\cdot}^{\mathfrak{g}}\,(\mathfrak{g}\in\{\mathfrak{su},\mathfrak{so}\})$ has equal integrals with respect to $\mu$ and $\nu$, then $\mu=\nu$.
Proposition 5.5 does not cover the case $\mathfrak{g}=\bigoplus_{k\geq 1}\mathfrak{t}(k;\mathbb{R})$ since it is not an algebra.
Now we specialise to the expected signature of a path. If the ROC $\rho(\mathbb{E}(S(\gamma)))=\infty$, then
$\mathbb{E}\big{[}\mathscr{D}_{\gamma}(M)\big{]}=\mathbb{E}\left[\widetilde{M}(S(\gamma))\right]=\widetilde{M}\big{[}\mathbb{E}(S(\gamma))\big{]}.$
By Proposition 5.5, the expected signature uniquely determines Radon measures on $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$ with $1\leq p<2$. Also, in view of § 1.3, $\widetilde{M}\big{[}\mathbb{E}(S(\gamma))\big{]}$ is the generating function of the expected signature. Hence, the expected signature can be recovered from
$\left\{\mathbb{E}\big{[}\mathscr{D}_{\gamma}(M)\big{]}:M\in\bigoplus_{k\geq 1}\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)\right\}$
by the same algorithm in Theorem 1.9. We thus arrive at the following:
Proposition 5.6.
Let $\mu$ and $\nu$ be Radon measures on $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$; $1\leq p<2$. Assume that their expected signatures exist and are of infinite ROC. The following are equivalent:
  • $\mu=\nu$;
  • $\mathbb{E}_{\gamma\sim\mu}\big{[}\mathscr{D}_{\gamma}(M)\big{]}=\mathbb{E}_{\gamma^{\prime}\sim\nu}\big{[}\mathscr{D}_{\gamma^{\prime}}(M)\big{]}$ for all $M\in\bigoplus_{k\geq 1}{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$.
Under the condition that the signature transform
$S:\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)\longrightarrow\mathcal{S}_{p}:=S\left(\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)\right)\subset\mathcal{G}\left(\mathbb{R}^{d}\right)$
is a homeomorphism onto its image, one may easily obtain an analogous result to Proposition 5.5 on the determination of Radon measures on $\mathcal{S}_{p}$.
Remark 5.7.
In Proposition 5.6, if one only considers measures on the time augmentation of the path space $\mathcal{V}_{0,p}\subset\mathcal{V}_{p}$, then every finite Borel measure is Radon as $\mathcal{V}_{0,p}$ is a Polish space (cf. [19, Proposition 5.36]). $\mathcal{V}_{0,p}$ is the closure of smooth paths under the $p$-variation norm, and is known as the space of geometric $p$-rough paths. See Friz–Victoir [19, Definition 9.15].
6. Numerical experiments on hypothesis testing
Now we present various numerical results based on the theories developed earlier in this paper, especially §5. In §6.1, we propose a new distance function for measures on path spaces. Then, in §6.2, we apply this proposed distance function on the task of hypothesis testing (i.e., deciding whether two stochastic processes have the same distribution on the path space). The numerical results demonstrate that our proposed method significantly outperforms the signature-based MMD [28], and achieves a comparable performance to PCFD-based approaches [16], while offering reduced time complexity by an order of magnitude.
For readers interested in reproducing the numerical experiments, the implementation code is available in the GitHub repository: https://github.com/DeepIntoStreams/UnitaryDevInversion.
6.1. A new distance of stochastic processes
Let $1\leq p<2$. In the sequel, we consider probability measures on $\tilde{\mathcal{V}}_{0,p}$, i.e., the time augmentation of the space of geometric $p$-rough paths, which is a Polish space (Remark 5.7). Denote
$\mathcal{P}_{\infty}\left(\tilde{\mathcal{V}}_{0,p}\right)$ $:=\bigg{\{}\text{Borel probability measures on $\tilde{\mathcal{V}}_{0,p}$ }$
 $\qquad\qquad\text{ whose expected signatures exist and with ${\rm ROC}=\infty$}\bigg{\}}.$
Then, by Proposition 5.6 and Remark 5.7 cf. [16]), the Path Characteristic Function (PCF)
$\Phi_{\mu}(M):=\mathbb{E}_{\gamma\sim\mu}\big{[}\mathscr{D}_{\gamma}(M)\big{]}=\mathbb{E}_{\gamma\sim\mu}\left[\widetilde{M}(S(\gamma))\right]$
fully characterises $\mu\in\mathcal{P}_{\infty}\left(\tilde{\mathcal{V}}_{0,p}\right)$, as long as its domain of definition contains $\bigoplus_{k\geq 1}{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$. This motivates us to introduce the following
@@ 6.1. A new distance of stochastic processes
Let $\mathcal{M}:=(\mathcal{M}_{k})_{k\geq 1}$, where $\mathcal{M}_{k}$ is a probability measure on $\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$. For any $\mu,\nu\in\mathcal{P}_{\infty}\left(\tilde{\mathcal{V}}_{0,p}\right)$, their Restricted Path Characteristic Function Distance (RPCFD) with respect to $\mathcal{M}$ is defined by
[…]
where $||\cdot||_{\rm HS}$ denotes the matrix Hilbert–Schmidt norm.
In view of Proposition 5.6, $\rm RPCFD$$_{\mathcal{M}}$ is a true distance as long as each $\mathcal{M}_{k}$ has a full support on $\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$. It implies that for any two different measures $\mu$ and $\nu$, there exists some $k\geq 1$ such that the $k$-th grade of ${\rm RPCFD}_{\mathcal{M}}^{2}(\mu,\nu)$ is non-zero; i.e.,
[…]
Since $\mathfrak{t}(1,\mathbb{R})\hookrightarrow\mathfrak{t}(2,\mathbb{R})\hookrightarrow\cdots\hookrightarrow\mathfrak{t}(k,\mathbb{R})\hookrightarrow\mathfrak{t}(k+1,\mathbb{R})\cdots$, one may work with a sufficiently large grade $k$ and the corresponding expectation $\mathbb{E}_{M\sim\mathcal{M}_{k}}$ in numerical experiments. However, larger values of $k$ lead to an increased computational complexity and a higher risk of overfitting. Therefore, cross-validation is employed in practice to select the most appropriate value of $k$.
Remark 6.2.
As $\mathfrak{t}(k,\mathbb{R})\subset{\mathfrak{su}}(k)$, RPCFD can be viewed as a restricted version of PCFD [16] without losing the characteristicity of measures on the path space. Since
@@
RPCFD has a significant dimension reduction effect over PCFD. Likewise, based on Proposition 5.5, we define the Orthogonal Path Characteristic Function Distance (OPCFD) by restricting the domain of $M$ to $\mathcal{L}\left(\mathbb{R}^{d};\bigoplus_{k=1}^{\infty}{\mathfrak{so}}(k)\right)$. Notice that the characteristicity of OPCFD does not require the existence of expected signature. As $dim({\mathfrak{so}}(k))=k(k-1)/2$, RPCFD reduces the dimension more significantly than OPCFD, so we focus more on RPCFD. See Appendix B for numerical experiments on OPCFD.
@@
Typically, empirical time series data are mostly observed at discrete time points. As in [19, Theorem 8.22 (i.3)], we approximate paths in $\tilde{\mathcal{V}}_{0,p}$ by interpolations (e.g., linear interpolation), and the corresponding Borel probability measures on $\tilde{\mathcal{V}}_{0,p}$ converge to the correct limit as the time step size tends to zero.
To enhance the discriminative power of RPCFD on given path distributions, given the grade $k\geq 1$, we propose the following trainable distance:
$\sup_{\mathcal{M}\in\mathcal{P}(\mathcal{L}(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})))}\text{RPCFD}^{2}_{\mathcal{M}}(\mu,\nu).$
@@
$\mathcal{M}\approx\frac{1}{K}\sum_{i=1}^{K}\delta_{M_{i}},$
where $M_{i}\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$, $i\in\{1,\cdots,K\}$, for some $K\geq 1$. Note that the linear maps $\{M_{i}\}_{1\leq i\leq K}$ can be fully parameterised. Therefore, the task of learning the optimal distribution $\mathcal{M}$ in (38) is transformed to learning the optimal parameters in the linear maps $\{M_{i}\}_{1\leq i\leq K}$ which attain the supremum of the “empirical RPCFD”:
[…]